# Decentralized COVID-19 Health System Leveraging Blockchain

### v1

@@ 1. Introduction
In 2008, Bencong developed the Bitcoin system(Nakamoto,, 2008). With the rising price of Bitcoin, the underlying technology foundation behind the system, blockchain technology, has gradually entered everyone’s vision, and has gradually become a research hotspot in the academic community. Today, blockchain layout frequently appears in government decision documents. The application of blockchain technology has become a hot research direction at present and in the future. More personnel have been devoted to research on blockchain technology. The development of blockchain based application systems is related to the development of the Internet field and the future economic development of countries and regions (Niu et al.,, 2024).
@@
With the vigorous development of Internet technology and applications, all walks of life have become inseparable from the Internet, including the medical industry. From the early queuing for registration and consultation, waiting for paper results, to the current simple online booking and inquiry, the Internet has greatly facilitated the medical industry. While enjoying convenience, one can also discover many problems, among which it is worth noting the storage and use of patient medical record information. According to rough statistics and calculations, the total storage capacity of medical data generated so far has reached an astonishing EB level. Especially in the past decade, the popularization of digitization has led to exponential growth of medical data, and the rapid development of medical informatization has also posed new challenges to medical information sharing(Dennis et al.,, 2016).
@@
Nowadays, more and more hospitals choose to store medical record information in the form of the Internet. Electronic medical records are gradually replacing traditional paper medical records, which greatly facilitates the development of the medical industry. When a patient needs to see a doctor, he does not need to provide his own paper medical records, but only need to query the personal medical record information in the electronic medical record management system. Through electronic medical records (EHR), doctors can easily understand the patient’s past illnesses and medical information, and can make a more detailed and comprehensive judgment of the condition. The treatment efficiency and treatment effect achieved by relying on electronic medical records are unmatched by traditional medical records. Meanwhile, when treating severe infectious diseases, EHR can share data with other excellent teams to help patients receive more comprehensive and rapid treatment(Zhu et al.,, 2024).
@@
However, due to the fact that traditional database systems are centralized systems and system data is fully controlled by relevant institutions, it means that data between different hospitals is not shared or is difficult to share for the medical and health system. In the actual operation process, when patients come to different medical institutions for treatment, they usually need to handle the sharing of medical records themselves. This goes back to the traditional solution, which is not what a complete electronic medical record management system hopes for. In addition, since electronic medical records are ultimately stored in medical institutions, the management and use of medical records are opaque to patients, which is also something we do not want to see. EHR stores patients’ personal information and medical records. Once attacked, it can lead to the leakage of sensitive information, such as patient privacy, causing security risks and doctor-patient conflicts(Zhong et al.,, 2023).
@@
With the development of information technology, it is inevitable that EHR will gradually replace traditional case management. Past data (Henry et al.,, 2016) also shows that various types of hospitals are increasing the proportion of electronic storage of cases. Therefore, solving the above problems has a very important practical application value. As the core technology of Bitcoin, blockchain has the properties of decentralization, data immutability, traceability, unforgeability, programmability, etc. The use of blockchain technology can establish a large-scale decentralized information management system for data sharing among multiple hospitals, which can facilitate hospitals to quickly query the medical records of various types of patients and solve the problem of medical data silos between hospitals. In addition, by designing a blockchain system, patients can have greater control over the use and information rights of their own medical records, providing better protection for their private data, such as medical records. And due to the tamper-proof and traceable nature of blockchain data, it can ensure that patient information will not be maliciously tampered with. Thus, build a decentralized, secure, and tamper proof medical information system.
@@ 2.1. Cryptography Technology
As a representative of blockchain, the Bitcoin system has not experienced any effective attacks at the system level since its inception, and its system security is evident. Among them, the security of the blockchain data layer mainly relies on cryptographic-related technologies, including hash algorithms, digital signatures, and encryption algorithms(Li et al.,, 2025). Hash algorithm is a one-way function mapping that outputs a fixed-length hash value for an input string of any length. The formula is shown by eq. (1).
@@
A large number of hash algorithms have been proposed so far, among which MD5 (Rivest,, 1992) and SHA (Eastlake 3rd and Jones,, 2001) are widely used. Among them, MD5 is widely used for data encryption and protection due to its universal, stable, and fast algorithm. However, with the continuous development of cryptanalysis, MD5 has been discovered to construct collision methods, which have been identified as insecure hash algorithms (Wang et al.,, 2005). In addition, the SHA algorithm (SHA1) with a shorter length y has also been cracked. The widely used SHA algorithm today is SHA2, including SHA256, SHA512, etc.
@@
The Bitcoin system uses SHA256 as the hash algorithm, and internal transactions in blocks adopt the Merkle Tree data structure, as shown in Fig 1. For each transaction data block, it serves as the input of the hash function, calculates its hash result, and uses it as input to the upper layer block. This process is repeated until a root hash is generated and stored in the block header. Therefore, based on the anti-collision characteristics of hash functions, it is not feasible to find a new transaction data block that keeps the hash result unchanged, thus ensuring that the data on the blockchain is tamper-proof(Kong et al.,, 2024).
@@
In addition, research has shown that hash functions are puzzle friendly, meaning that the overall computation is memoryless. This characteristic is applied in the mining process of blockchain systems built on the POW mechanism, ensuring that miners’ ability to obtain accounting rights is proportional to the computing power of nodes, and nodes with stronger computing power will not gain disproportionate advantages (Liu et al., 2024b, ).
@@
A digital signature (also known as a public key digital signature) is a string of digits that cannot be forged by others and can only be generated by the sender of the information. This string is also an effective proof of the authenticity of the information sent by the sender. In 1977, the famous RSA algorithm was proposed, and modern cryptography developed the most important public key cryptography system. It can be said that without the public key cryptography system represented by RSA, there would be no digital signature. To implement digital signatures, the following requirements must first be met :
  • Authentication.Considering traditional symmetric encryption algorithms, encryption and decryption correspond to the signature and verification of digital signatures. However, since the private key owner is a single individual, it cannot meet the application scenarios of digital signatures. The emergence of public key cryptography allows signers to encrypt signatures with their own private keys, while validators can use publicly available signer public keys for decryption and verification, effectively verifying the signer’s identity.
  • Non-Repudiation.After the signer signs the content, the characteristics of the signature algorithm and public key information cannot infer the private key data, which means that only the signer knows the private key. For a signature that can be successfully verified by the public key, the signer cannot deny it.
  • Data Integrity.Due to the correlation between the input of the signature algorithm and the signed data file, this correlation can be achieved by directly using the file’s data as input, or by using the file’s summary information as input for operational efficiency. This means that the information of the signed file cannot be modified.
Taking Bitcoin as an example, the Bitcoin system uses the elliptic curve-based digital signature algorithm ECDSA. When a node publishes a transfer message, it first needs to sign the transaction information with a private key. After receiving the transaction information, the miner node verifies the transaction data signature with the signer’s public key. The characteristics of digital signature algorithms ensure the verifiability, integrity, and non-repudiation of transaction data.
@@ 2.2. Consensus Mechanism
As blockchain is a decentralized system, for public chain systems, each system node can independently choose to join this distributed system. Nodes communicate with each other through a P2P network, and each node is jointly responsible for the normal operation of the blockchain system. For such a complex distributed network system, in order to enable nodes to reach a consensus and jointly build the blockchain system they belong to, the design of the blockchain consensus algorithm is extremely crucial (Bu et al., 2025a, ).
@@
Since the concept of blockchain was proposed, consensus algorithms have always been a key research direction in blockchain technology. Among them, typical consensus algorithms include the Proof of Work (POW) mechanism, the Proof of Stake (POS) mechanism, the Delegated Proof of Stake (DPOS) mechanism (Larimer,, 2014), and the Practical Byzantine Fault Tolerance (PBFT) algorithm(Castro et al.,, 1999).
@@
The PoW algorithm originates from Bitcoin. The core idea of this algorithm is to allow nodes to compete for the right to record transactions by competing in computing power. To make nodes willing to serve the blockchain system, the system sets up corresponding incentive mechanisms, rewarding nodes that obtain the right to record transactions with a certain amount of tokens. In the Bitcoin system, it is stipulated that the hash value of the block header must be less than a set value. Due to the characteristics of the hash algorithm, nodes can only continuously modify the random number and keep trying to find a block whose hash value is less than the difficulty value, and then broadcast the corresponding block for verification by other nodes.
@@
The PoS algorithm was proposed based on the PoW algorithm. In response to the huge energy consumption problem caused by the computing power competition in the PoW algorithm’s mining process, PoS adopts the approach of granting the right to record transactions to the node with the highest stake in the system. This stake can be reflected in the blockchain system of cryptocurrencies as the quantity and holding time of tokens held by the node (Liu and Li,, 2025). This means that nodes with a larger number of tokens and longer holding periods have a higher possibility of obtaining the right to record transactions. The DPoS algorithm is a further development of the PoS algorithm. Its core idea is to use a voting election method to select nodes with the right to record transactions. First, N representative nodes need to be elected, and these representative nodes take turns to obtain the right to record transactions. According to the actual operation, representative nodes can be re-elected.
@@
The BFT algorithm(Lamport et al.,, 2019) takes into account the possibility of malicious nodes in the system. When the number of Byzantine nodes is less than one-third of the total number of nodes, nodes can reach a consensus normally. Representative algorithms of the BFT algorithm include the PBFT algorithm, which mainly optimizes the efficiency of the BFT. The PBFT algorithm includes three stages: pre-preparation, preparation, and commit. Since nodes need to communicate with each other, the communication complexity of the system increases with an increasing number of nodes, and the communication cost is relatively high. Therefore, this algorithm is not suitable for public chain scenarios with a large number of nodes, but is suitable for consortium chains and private chains.
@@ 2.3. Smart Contract
In 1997, cryptographer Nick Szabo first proposed the concept of smart contracts(Zou et al.,, 2025), where Szabo hoped to eliminate intermediaries and make contracts automatically effective as long as certain conditions were met.
The emergence of blockchain technology has driven the development of smart contracts. No need for a third party, the contract is a computer protocol that exists in code form. Its essence is a computer program written by programmers and stored in blocks, using deterministic algorithms and data sources, and satisfying terminability. When the network state changes and meets the initial conditions set, the ’miner’ will automatically execute the smart contract content. In Ethereum, smart contracts are considered programs that can directly control digital assets. Ethereum and Hyperledger Fabric are two representative technology platforms for applying smart contracts (Li et al.,, 2021).
@@ 2.4. Searchable Encryption Technology
When storing information on an untrusted server, such as email servers and file servers, we hope to protect the stored data in an encrypted form to reduce security and privacy risks. But such measures taken for safety will cause us to lose some of our original functions. For example, if a client wants to retrieve its stored files through keywords, ordinary encrypted data algorithms cannot match the ciphertext with the given search term, making it impossible to achieve the corresponding search function. Therefore, in the past, it was not known how to make the data storage server perform searches and answer queries without losing data confidentiality (Li et al.,, 2017).
@@ 2.5. Access Control Technology for Permission
The concept of proxy re-encryption was first proposed by BLAZE et al.(Blaze et al.,, 1998), and its core idea is to use a proxy to perform secondary encryption on stored encrypted data. The result of the secondary encryption can be decrypted by the private key of the data’s sharer, and the proxy cannot obtain any plaintext information during the process (Lang et al.,, 2014).
@@ 3.1.2. Running Process
    • Key Generation: Based on the requirements of DO and DU, CA can generate the keys needed for DO and DU, and manage the entire key usage process through the organization’s principles of master key, secondary key, and tertiary key. In addition, considering the possible issue of users’ lack of trust in the security of CA management keys, support for user clients to generate keys locally, provide public key information registration, and subsequently identify identity through certificates (Li et al., 2024c, ).
@@ 3.2.1. Universal Module
The general function module includes various general functions of system users, which have similar implementation processes. Considering the convenience of later maintenance and program scalability, functions with similar functions or implementation processes can be grouped into one module(Ting et al.,, 2017).
@@ 3.2.2. Data Owner Module
COVID-19 health data belongs to the privacy of DO. The design of this system needs to ensure that DO has absolute control over the data. As stipulated at the beginning of this chapter, all data is divided into public data and private data according to whether the information is encrypted or not. For public data, the access control mechanism of the system is set up, which is not public to DU. When DU wants to access public data, it also needs to apply for access permission from DO(Bu et al., 2025b, ).

### v2

@@ 1. Introduction
In 2008, Bencong developed the Bitcoin system(Nakamoto, 2008). With the rising price of Bitcoin, the underlying technology foundation behind the system, blockchain technology, has gradually entered everyone’s vision(Panda et al., 2023), and has gradually become a research hotspot in the academic community. Today, blockchain layout frequently appears in government decision documents. The application of blockchain technology has become a hot research direction at present and in the future. More personnel have been devoted to research on blockchain technology (Dong et al., 2023). The development of blockchain based application systems is related to the development of the Internet field and the future economic development of countries and regions.
@@
With the vigorous development of Internet technology and applications, all walks of life have become inseparable from the Internet, including the medical industry (Li, 2023; Hong, 2022; Zhou and Sheng, 2022). From the early queuing for registration and consultation, waiting for paper results, to the current simple online booking and inquiry, the Internet has greatly facilitated the medical industry. While enjoying convenience, one can also discover many problems, among which it is worth noting the storage and use of patient medical record information. According to rough statistics and calculations, the total storage capacity of medical data generated so far has reached an astonishing EB level (Mahajan and Junnarkar, 2023). Especially in the past decade, the popularization of digitization has led to exponential growth of medical data, and the rapid development of medical informatization has also posed new challenges to medical information sharing(Dennis et al., 2016).
@@
Nowadays, more and more hospitals choose to store medical record information in the form of the Internet. Electronic medical records are gradually replacing traditional paper medical records, which greatly facilitates the development of the medical industry. When a patient needs to see a doctor, he does not need to provide his own paper medical records, but only need to query the personal medical record information in the electronic medical record management system. Through electronic medical records (EHR), doctors can easily understand the patient’s past illnesses and medical information, and can make a more detailed and comprehensive judgment of the condition (Murala et al., 2023). The treatment efficiency and treatment effect achieved by relying on electronic medical records are unmatched by traditional medical records. Meanwhile, when treating severe infectious diseases, EHR can share data with other excellent teams to help patients receive more comprehensive and rapid treatment (Zhu et al., 2024).
@@
However, due to the fact that traditional database systems are centralized systems and system data is fully controlled by relevant institutions (Kassen, 2022; Schlatt et al., 2022; Tyagi, 2023), it means that data between different hospitals is not shared or is difficult to share for the medical and health system. In the actual operation process, when patients come to different medical institutions for treatment, they usually need to handle the sharing of medical records themselves. This goes back to the traditional solution, which is not what a complete electronic medical record management system hopes for. In addition, since electronic medical records are ultimately stored in medical institutions, the management and use of medical records are opaque to patients, which is also something we do not want to see. EHR stores patients’ personal information and medical records. Once attacked, it can lead to the leakage of sensitive information, such as patient privacy, causing security risks and doctor-patient conflicts(Zhong et al., 2023).
@@
With the development of information technology, it is inevitable that EHR will gradually replace traditional case management. Past data (Henry et al., 2016) also shows that various types of hospitals are increasing the proportion of electronic storage of cases. Therefore, solving the above problems has a very important practical application value. As the core technology of Bitcoin, blockchain has the properties of decentralization, data immutability, traceability, unforgeability, programmability, etc (Gharavi et al., 2024). The use of blockchain technology can establish a large-scale decentralized information management system for data sharing among multiple hospitals, which can facilitate hospitals to quickly query the medical records of various types of patients and solve the problem of medical data silos between hospitals. In addition, by designing a blockchain system, patients can have greater control over the use and information rights of their own medical records, providing better protection for their private data, such as medical records. And due to the tamper-proof and traceable nature of blockchain data, it can ensure that patient information will not be maliciously tampered with (Zhang et al., 2022). Thus, build a decentralized, secure, and tamper proof medical information system.
@@ 2.1. Cryptography Technology
As a representative of blockchain, the Bitcoin system has not experienced any effective attacks at the system level since its inception, and its system security is evident. Among them, the security of the blockchain data layer mainly relies on cryptographic-related technologies, including hash algorithms, digital signatures, and encryption algorithms. Hash algorithm is a one-way function mapping that outputs a fixed-length hash value for an input string of any length. The formula is shown by eq. (1).
@@
A large number of hash algorithms have been proposed so far, among which MD5 (Rivest, 1992) and SHA (Eastlake 3rd and Jones, 2001) are widely used. Among them, MD5 is widely used for data encryption and protection due to its universal, stable, and fast algorithm. However, with the continuous development of cryptanalysis, MD5 has been discovered to construct collision methods, which have been identified as insecure hash algorithms (Wang et al., 2005). In addition, the SHA algorithm (SHA1) with a shorter length y has also been cracked. The widely used SHA algorithm today is SHA2, including SHA256, SHA512, etc.
@@
The Bitcoin system uses SHA256 as the hash algorithm, and internal transactions in blocks adopt the Merkle Tree data structure, as shown in Fig 1. For each transaction data block, it serves as the input of the hash function, calculates its hash result, and uses it as input to the upper layer block. This process is repeated until a root hash is generated and stored in the block header. Therefore, based on the anti-collision characteristics of hash functions, it is not feasible to find a new transaction data block that keeps the hash result unchanged, thus ensuring that the data on the blockchain is tamper-proof(Kong et al., 2024).
@@
In addition, research has shown that hash functions are puzzle friendly, meaning that the overall computation is memoryless. This characteristic is applied in the mining process of blockchain systems built on the POW mechanism, ensuring that miners’ ability to obtain accounting rights is proportional to the computing power of nodes, and nodes with stronger computing power will not gain disproportionate advantages.
@@
A digital signature (also known as a public key digital signature) is a string of digits that cannot be forged by others and can only be generated by the sender of the information. This string is also an effective proof of the authenticity of the information sent by the sender. In 1977, the famous RSA algorithm was proposed, and modern cryptography developed the most important public key cryptography system (Imam et al., 2022). It can be said that without the public key cryptography system represented by RSA, there would be no digital signature. To implement digital signatures, the following requirements must first be met :
  • Authentication.Considering traditional symmetric encryption algorithms, encryption and decryption correspond to the signature and verification of digital signatures. However, since the private key owner is a single individual (Halak et al., 2022), it cannot meet the application scenarios of digital signatures. The emergence of public key cryptography allows signers to encrypt signatures with their own private keys, while validators can use publicly available signer public keys for decryption and verification, effectively verifying the signer’s identity.
  • Non-Repudiation.After the signer signs the content, the characteristics of the signature algorithm and public key information cannot infer the private key data, which means that only the signer knows the private key (Liu et al., 2023). For a signature that can be successfully verified by the public key, the signer cannot deny it.
  • Data Integrity.Due to the correlation between the input of the signature algorithm and the signed data file, this correlation can be achieved by directly using the file’s data as input, or by using the file’s summary information as input for operational efficiency (Hong et al., 2023). This means that the information of the signed file cannot be modified.
Taking Bitcoin as an example, the Bitcoin system uses the elliptic curve-based digital signature algorithm ECDSA (Ulla and Sakkari, 2025; Huynh et al., 2023; Gani et al., 2024). When a node publishes a transfer message, it first needs to sign the transaction information with a private key. After receiving the transaction information, the miner node verifies the transaction data signature with the signer’s public key. The characteristics of digital signature algorithms ensure the verifiability, integrity, and non-repudiation of transaction data.
@@ 2.2. Consensus Mechanism
As blockchain is a decentralized system, for public chain systems, each system node can independently choose to join this distributed system. Nodes communicate with each other through a P2P network, and each node is jointly responsible for the normal operation of the blockchain system. For such a complex distributed network system, in order to enable nodes to reach a consensus and jointly build the blockchain system they belong to, the design of the blockchain consensus algorithm is extremely crucial (Bu et al., 2025).
@@
Since the concept of blockchain was proposed, consensus algorithms have always been a key research direction in blockchain technology. Among them, typical consensus algorithms include the Proof of Work (POW) mechanism, the Proof of Stake (POS) mechanism, the Delegated Proof of Stake (DPOS) mechanism (Larimer, 2014), and the Practical Byzantine Fault Tolerance (PBFT) algorithm(Castro et al., 1999).
@@
The PoW algorithm originates from Bitcoin. The core idea of this algorithm is to allow nodes to compete for the right to record transactions by competing in computing power. To make nodes willing to serve the blockchain system, the system sets up corresponding incentive mechanisms, rewarding nodes that obtain the right to record transactions with a certain amount of tokens. In the Bitcoin system, it is stipulated that the hash value of the block header must be less than a set value. Due to the characteristics of the hash algorithm, nodes can only continuously modify the random number and keep trying to find a block whose hash value is less than the difficulty value (Ali et al., 2023), and then broadcast the corresponding block for verification by other nodes.
@@
The PoS algorithm was proposed based on the PoW algorithm. In response to the huge energy consumption problem caused by the computing power competition in the PoW algorithm’s mining process, PoS adopts the approach of granting the right to record transactions to the node with the highest stake in the system (Rahman et al., 2024; Fior, 2024; Feng et al., 2022). This stake can be reflected in the blockchain system of cryptocurrencies as the quantity and holding time of tokens held by the node. This means that nodes with a larger number of tokens and longer holding periods have a higher possibility of obtaining the right to record transactions (Mišić et al., 2023). The DPoS algorithm is a further development of the PoS algorithm (Selvam et al., 2024). Its core idea is to use a voting election method to select nodes with the right to record transactions. First, N representative nodes need to be elected, and these representative nodes take turns to obtain the right to record transactions. According to the actual operation, representative nodes can be re-elected.
@@
The BFT algorithm(Lamport et al., 2019) takes into account the possibility of malicious nodes in the system. When the number of Byzantine nodes is less than one-third of the total number of nodes, nodes can reach a consensus normally. Representative algorithms of the BFT algorithm include the PBFT algorithm, which mainly optimizes the efficiency of the BFT. The PBFT algorithm includes three stages: pre-preparation, preparation, and commit. Since nodes need to communicate with each other, the communication complexity of the system increases with an increasing number of nodes, and the communication cost is relatively high. Therefore, this algorithm is not suitable for public chain scenarios with a large number of nodes, but is suitable for consortium chains and private chains.
@@ 2.3. Smart Contract
In 1997, cryptographer Nick Szabo first proposed the concept of smart contracts(Zou et al., 2025), where Szabo hoped to eliminate intermediaries and make contracts automatically effective as long as certain conditions were met.
The emergence of blockchain technology has driven the development of smart contracts. No need for a third party, the contract is a computer protocol that exists in code form (Sharma et al., 2023; Afraz et al., 2023; Ahmed et al., 2024). Its essence is a computer program written by programmers and stored in blocks, using deterministic algorithms and data sources, and satisfying terminability. When the network state changes and meets the initial conditions set, the ’miner’ will automatically execute the smart contract content. In Ethereum, smart contracts are considered programs that can directly control digital assets. Ethereum and Hyperledger Fabric are two representative technology platforms for applying smart contracts.
@@ 2.4. Searchable Encryption Technology
When storing information on an untrusted server, such as email servers and file servers, we hope to protect the stored data in an encrypted form to reduce security and privacy risks (Gupta et al., 2022). But such measures taken for safety will cause us to lose some of our original functions. For example, if a client wants to retrieve its stored files through keywords, ordinary encrypted data algorithms cannot match the ciphertext with the given search term, making it impossible to achieve the corresponding search function. Therefore, in the past, it was not known how to make the data storage server perform searches and answer queries without losing data confidentiality (Wu et al., 2022).
@@ 2.5. Access Control Technology for Permission
The concept of proxy re-encryption was first proposed by BLAZE et al.(Blaze et al., 1998), and its core idea is to use a proxy to perform secondary encryption on stored encrypted data. The result of the secondary encryption can be decrypted by the private key of the data’s sharer, and the proxy cannot obtain any plaintext information during the process (Lang et al., 2014).
@@ 3.1.2. Running Process
    • Key Generation: Based on the requirements of DO and DU, CA can generate the keys needed for DO and DU, and manage the entire key usage process through the organization’s principles of master key, secondary key, and tertiary key. In addition, considering the possible issue of users’ lack of trust in the security of CA management keys, support for user clients to generate keys locally, provide public key information registration, and subsequently identify identity through certificates (Li et al., 2024c).
@@ 3.2.1. Universal Module
The general function module includes various general functions of system users, which have similar implementation processes. Considering the convenience of later maintenance and program scalability, functions with similar functions or implementation processes can be grouped into one module(Ting et al., 2017).
@@ 3.2.2. Data Owner Module
COVID-19 health data belongs to the privacy of DO. The design of this system needs to ensure that DO has absolute control over the data. As stipulated at the beginning of this chapter, all data is divided into public data and private data according to whether the information is encrypted or not. For public data, the access control mechanism of the system is set up, which is not public to DU. When DU wants to access public data, it also needs to apply for access permission from DO.